A local bagel shop produces bagels (B) and croissants (C). Each bagel requires 6 ounces of flour, 1 gram of yeast, and 2 tablespoons of sugar. A croissant requires 3 ounces of flour, 1 gram of yeast, and 4 tablespoons of sugar. The company has 6,600 ounces of flour, 1,400 grams of yeast, and 4,800 tablespoons of sugar available for today’s baking. Bagel profits are 20 cents each and croissant profits are 30 cents each. The shop wishes to maximize profits. Which of the following is not a feasible solution?

Multiple Choice (

B, C) = (1100, 0)

(B, C) = (800, 600)

(B, C) = (0, 1100)

(B, C) = (0, 1400)
Answer #1

The constain equations will be

By flour

6B+3C<=6600

yeast

B+C <= 1400

sugar

2B+4C <= 4800

Profit = 0.20 B + 0.30C

(B, C) = (0, 1400) is not in the fesible region